Consider a charge (+5.0μC) fixed at a site with another charge (charge +3.0μC, mass 6.0μg) moving in the neighboring space. (a) Evaluate the potential energy of when it is 4.0 cm from . (b) If starts from rest from a point 4.0 cm from , what will be its speed when it is 8.0 cm from ? (Note:is held fixed in its place.




1
Expert's answer
2022-02-10T13:38:10-0500

"q_1=+5\mu C=5*10^{-6}C"

"q_2=+3\mu C=+3*10^{-6}C"

"m = 6\mu g=6*10^{-9}kg"

"r=4cm=0.4m"

"W-?"

"W=k*\large\frac{q_1*q_2}{r}=\frac{9*10^9*5*10^{-6}*3*10^{-6}}{0.4}=0.3375J"